# One Perturbation is Enough: On Generating Universal Adversarial Perturbations against Vision-Language Pre-training Models

### v2

@@ 4 Evaluation
| Flickr30k | ALBEF | GAP | 69.78 | 81.59 | 22.15 | 29.97 | 6.61 | 18.37 | 23.4 | 37.54 | 29.92 | 44.29 | 16.09 | 28.12 |
| Ours | 90.13 | 88.82 | 62.11 | 64.48 | 20.53 | 39.38 | 43.1 | 65.93 | 54.4 | 72.51 | 44.79 | 56.36 | ||
| TCL | GAP | 33.5 | 40.61 | 82.41 | 80.67 | 6.61 | 17.79 | 21.55 | 38.56 | 30.57 | 45.48 | 21.45 | 31.82 | |
| Ours | 50.26 | 56.29 | 94.93 | 90.64 | 14.94 | 33.96 | 46.92 | 66.41 | 52.98 | 70.66 | 35.75 | 52.52 | ||
| X-VLM | GAP | 16.14 | 24.43 | 17.08 | 26.2 | 90.24 | 85.98 | 24.51 | 41.15 | 42.62 | 53.08 | 16.19 | 25.74 | |
| Ours | 24.46 | 47.77 | 29.19 | 50.15 | 93.29 | 91.9 | 43.47 | 66.03 | 59.2 | 72.79 | 32.39 | 52.24 | ||
| CLIP | GAP | 11.72 | 23.34 | 15.32 | 26.39 | 8.54 | 20.48 | 85.73 | 90.45 | 48.83 | 60.78 | 14.83 | 26.46 | |
| Ours | 23.23 | 38.67 | 25.05 | 41.79 | 15.85 | 35.59 | 88.92 | 93.05 | 66.06 | 75.42 | 26.71 | 45.7 | ||
| CLIP | GAP | 13.57 | 25.21 | 19.05 | 28.87 | 11.59 | 23.13 | 27.46 | 43.16 | 73.18 | 81.6 | 15.25 | 27.94 | |
| Ours | 15.31 | 38.93 | 19.77 | 43.72 | 17.17 | 41.65 | 39.9 | 64.82 | 81.74 | 88.9 | 22.19 | 46.11 | ||
| BLIP | GAP | 12.23 | 23.94 | 14.49 | 25.44 | 6.91 | 17.81 | 20.32 | 37 | 26.81 | 43.59 | 47.21 | 73.33 | |
| Ours | 32.17 | 44.4 | 33.44 | 44.51 | 18.6 | 35.53 | 43.35 | 60.26 | 48.96 | 66.95 | 71.82 | 82.82 | ||
| MSCOCO | ALBEF | GAP | 82.65 | 84.35 | 53.6 | 45.46 | 15.09 | 15.64 | 25.18 | 29.94 | 28.06 | 35.28 | 37.44 | 33.61 |
| Ours | 96.18 | 95.09 | 82.49 | 76.24 | 39.97 | 48.58 | 59.71 | 67.05 | 61.27 | 70.8 | 59.18 | 63.89 | ||
| TCL | GAP | 55.92 | 48.22 | 95.16 | 92.29 | 17.34 | 17.01 | 28.73 | 31.19 | 32.27 | 39.81 | 43.59 | 39.64 | |
| Ours | 76.62 | 71.17 | 96.72 | 93.88 | 42.99 | 48.4 | 70.32 | 79.08 | 74.1 | 82.97 | 62.35 | 66.97 | ||
| X-VLM | GAP | 26.35 | 23.72 | 27.8 | 22.91 | 95.1 | 88.84 | 32.39 | 38.16 | 52 | 55.4 | 24.67 | 22.65 | |
| Ours | 51.46 | 65.71 | 52.8 | 64.99 | 98.89 | 95.79 | 67.42 | 75.45 | 75.49 | 82.58 | 55.74 | 66.7 | ||
| CLIP | GAP | 35.96 | 31.91 | 37.33 | 32.56 | 33.42 | 29.25 | 97.71 | 96.04 | 74.63 | 74.67 | 33.47 | 31.99 | |
| Ours | 46.92 | 53.89 | 46.03 | 50.87 | 41.49 | 48.6 | 98.74 | 98.01 | 81.58 | 86.5 | 47.35 | 57.55 | ||
| CLIP | GAP | 28.67 | 27.51 | 29.84 | 27.69 | 26.4 | 24.81 | 39.64 | 40.53 | 90.34 | 91.56 | 24.99 | 26.18 | |
| Ours | 33.38 | 46.68 | 40.61 | 50.76 | 35.34 | 46.95 | 63.83 | 70.15 | 94.89 | 94.42 | 37.38 | 53.06 | ||
| BLIP | GAP | 35.55 | 38.75 | 35.62 | 33.79 | 22.7 | 21.25 | 32.05 | 35.8 | 40.93 | 45.58 | 73.46 | 72.37 | |
| Ours | 61.95 | 60.92 | 60.95 | 59.57 | 51.81 | 52.53 | 62.23 | 72.51 | 69.61 | 78.44 | 91.67 | 90.42 | ||
4.1 Experimental Setup
Downstream tasks and datasets. We conduct a comprehensive study of C-PGC on four downstream V+L tasks, including image-text retrieval (ITR), image captioning (IC), visual grounding (VG), and visual entailment (VE). For ITR tasks, we employ the Flickr30K (Plummer et al., 2015) and MSCOCO (Lin et al., 2014) datasets which are commonly used in previous works (Zhang et al., 2022; Lu et al., 2023). The MSCOCO is also adopted for evaluating the IC task and we test VG and VE tasks on SNLI-VE (Xie et al., 2019) and RefCOCO+ (Yu et al., 2016) respectively.
Surrogate models and victim models. We conduct experiments on a wide range of VLP models including ALBEF (Li et al., 2021), TCL (Yang et al., 2022), X-VLM (Zeng et al., 2022), CLIP (Radford et al., 2021), CLIP (Radford et al., 2021), and BLIP (Li et al., 2022). Note that for different V+L tasks, we correspondingly select different VLP models for evaluation based on their capability (Wang et al., 2024). For instance, among the six considered VLP models, only ALBEF, TCL, and X-VLM can handle VG tasks, while only ALBEF and TCL can deal with VE tasks.
Baselines. To better reveal the superiority of our proposed method in attacking VLP models, we transplant a representative and powerful algorithm GAP (Poursaeed et al., 2018) to the multimodal attack scenarios by appropriately modifying its original loss function (Lu et al., 2023).
Implementation details. Following (Lu et al., 2023), we adopt Karpathy split (Karpathy & Fei-Fei, 2015) to preprocess the dataset and build the test set for evaluation. The test set is disjoint with the generator’s training data for rigorous assessment. To ensure the perturbation invisibility, we follow (Wang et al., 2024) and limit the perturbation budgets to and to . During the augmentation, we resize the original images into five scales , and apply Gaussian noise with a mean of and a standard deviation of . See Appendix C for more details.
4.2 Universal Attack Effectiveness
To align with previous adversarial attacks on VLP models (Zhang et al., 2022; Lu et al., 2023), we first consider the typical V+L task of image-text retrieval and report the ASR computed based on R@1 retrieval results. Appendix 7 provides results of R@5 and R@10. Experimental results on MSCOCO and Flickr30K across six different VLP models are presented in Table 1. We also provide a visualization example of the image retrieval on the MSCOCO dataset in Figure 4.
White-box attack performance. By observing the white-box ASR in the gray-shaded area, we demonstrate that the proposed algorithm stably achieves excellent ASR on all the evaluated VLP models, validating the outstanding capability of the produced UAP. With only a single pair of perturbations, we reach a noteworthy average white-box ASR of over 90% on two large datasets in terms of both TR and IR tasks. Especially on the MSCOCO dataset, our method achieves over 95% average ASR on ITR tasks across six surrogate models. Compared with the GAP, the proposed method significantly improves the average fooling rates by nearly 10%, confirming the great validity of our suggested multimodal contrastive-learning mechanism. Essentially, the exceptional performance stems from the efficacy of our generated UAP in destroying the alignment between the image and text modalities, thereby misleading the VLP model during inference.
Black-box attack performance. We also conduct thorough experiments regarding the adversarial transferability of the generated UAP by transferring from surrogate models to other inaccessible models. As demonstrated in Table 1, the proposed C-PGC displays great attack performance in the more realistic black-box scenarios, e.g., 82.97% from TCL to CLIP on MSCOCO for IR tasks. We highlight that the advantage of C-PGC over GAP (Poursaeed et al., 2018) is greatly amplified in the more challenging black-box scenarios, which achieves a significant average improvement of 18.36% and 26.32% for Flickr30K and MSCOCO respectively. These experimental results indicate that our generative contrastive learning framework does not overly rely on the encoded feature space tailored to the surrogate model. Conversely, it is well capable of transferring to breaking the multimodal alignment of other unseen target models, thus attaining superior adversarial transferability.
@@ 4.2 Universal Attack Effectiveness
Defense Strategies. We next analyze several defense strategies to mitigate the potential harm brought by the proposed C-PGC. Specifically, we totally align with TMM (Wang et al., 2024) and consider several input preprocessing-based schemes, including image smoothing (Ding et al., 2019) (Gaussian, medium, average smoothing), JPEG compression (Dziugaite et al., 2016), NRP (Naseer et al., 2020), and the prevalent DiffPure (Nie et al., 2022), a powerful purification defense using diffusion models. For adversarial text correction, we choose the LanguageTool (LT) (Wang et al., 2024), which has been widely adopted in various scenarios due to its universality and effectiveness.
The attack results in Table 2 demonstrate that the proposed attack still attains great ASR against different powerful defenses. It also indicates that NRP+LT would be a decent choice to alleviate the threat brought by C-PGC. Another noteworthy finding is that, although DiffPure (Nie et al., 2022) exhibits remarkable performance in defending attacks in classification tasks, its ability is greatly reduced in V+L scenarios since the denoise process could also diminish some texture or semantic information that is critical for VLP models, thereby acquiring unsatisfactory defense effects.
4.3 Evaluation on More Downstream Tasks
We further demonstrate C-PGC’s ability to destroy the multimodal alignment by presenting more results on diverse V+L tasks. Specifically, we consider Image Captioning (IC), Visual Grounding (VG), and Visual Entailment (VE). The results of VE are shown in Appendix 7 due to space limit.
| Source | B@4 | METEOR | ROUBE_L | CIDEr | SPICE |
| Baseline | 39.7 | 31.0 | 60.0 | 133.3 | 23.8 |
| ALBEF | 30.1 | 23.7 | 51.2 | 92.5 | 17.5 |
| TCL | 29.5 | 23.5 | 51.0 | 88.9 | 17.3 |
| BLIP | 21.2 | 19.1 | 45.5 | 62.5 | 13.7 |
Image captioning. The objective of IC is to generate text descriptions relevant to the semantic content based on the given image. We use ALBEF, TCL, and BLIP as source models and attack the commonly used captioning model BLIP. Similar to SGA (Lu et al., 2023), several typical evaluation metrics of IC are calculated to measure the quality of generated captions, including BLEU (Papineni et al., 2002), METEOR (Banerjee & Lavie, 2005), ROUGE (Lin, 2004), CIDEr (Vedantam et al., 2015), and SPICE (Anderson et al., 2016). The results in Table 3 demonstrate that our algorithm again displays prominent attack effects, e.g., the crated UAP induces notable drops of 10.2% and 9% in the B@4 and ROUGE_L respectively when transferred from TCL to BLIP.
Target Baseline ALBEF TCL X-VLM Val TestA TestB Val TestA TestB Val TestA TestB Val TestA TestB ALBEF 58.4 65.9 46.2 37.1 39.8 32.0 42.2 46.9 35.2 37.6 40.2 33.0 TCL 59.6 66.8 48.1 43.6 47.8 36.9 39.0 41.4 33.6 39.5 41.7 34.1 X-VLM 70.8 67.8 61.8 51.8 54.7 47.7 52.7 55.9 47.8 33.1 34.7 28.8
Visual grounding. This is another common V+L task, which aims to locate the correct position in an image based on a given textual description. We conduct experiments on RefCOCO+ using ALBEF, TCL, and X-VLM as source and target models. Table 4 indicates that C-PGC brings a notable negative impact on the localization accuracy in both white-box and black-box settings, again verifying that the produced UAP strongly breaks the cross-modal interaction and alignment.
4.4 Ablation Study
This subsection employs the representative ALBEF (Li et al., 2021) model as the surrogate model to provide sufficient ablation studies on Flickr30K. We begin our analysis on the contribution of each proposed technique. Subsequently, we examine the sensitivity of certain hyperparameters.
The effect of and . To investigate the impact of the proposed loss terms, we introduce two variants C-PGCCL and C-PGCDis that remove and from the training loss respectively. As shown in Table 5, the removal of leads to significant degradation, particularly for black-box transferable attacks. e.g., a 27.12% ASR drop in TR tasks from ALBEF to TCL. This validates the considerable contribution of to guarantee a successful attack. Regarding the influence of , we demonstrate that this unimodal guidance can further enhance attacks on the basis of , e.g., a 10.59% increase in the ASR of TR tasks for white-box attacks on ALBEF. The proposed two loss terms complement each other and jointly underpin the generalizability of the generated UAP.
@@ 4.4 Ablation Study
| C-PGC | 90.13 | 88.82 | 62.11 | 64.48 | 20.53 | 39.38 | 43.1 | 65.93 | 54.4 | 72.51 | 44.79 | 56.36 |
| C-PGCDis | 79.54 | 82.46 | 56.52 | 62.21 | 20.24 | 38.26 | 39.78 | 65.1 | 52.2 | 71.01 | 42.43 | 55.52 |
The effect of positive sample selection. To validate the farthest selection strategy when constructing positive samples, we design another variant C-PGCRand that adopts randomly sampled data points as positive samples. Results in Table 5 reveal the necessity of the proposed farthest selection strategy as it brings an average improvement of 25.96% in white-box ASR and 4.95% in black-box ASR. Moreover, we can also conclude that if the positive samples are not adequately defined, adding would even severely harm the white-box performance (see C-PGCCL and C-PGCRand).
The effect of cross-modal conditions. As aforementioned, cross-attention (CA) modules are introduced into the generator to exploit cross-modal information. We then design C-PGCCA that cancels these CA layers to explore their influence. As expected, it causes a notable 9.78% average decrease across six target models, confirming the vital role of cross-modal knowledge. An interesting finding is that C-PGCCA induces a more pronounced drop in black-box attacks than white-box ones, indicating that cross-modal conditions exert a greater contribution to the adversarial transferability.
Different regulatory factor . The value of is a critical factor as it adjusts the scales of the two loss terms and . We explore the attack performance under various values of to confirm the optimal value. Figure 5 indicates that achieves superior performance.
Different perturbation budgets and . As shown in Figure 6, we analyze varying perturbation budgets for and . Generally, the ASR increases with the larger perturbation magnitudes. Note that when , C-PGC’s performance is severely compromised since the budget is too small to allow the UAP to carry enough information required to generalize to diverse data samples. We also find that the improvement slows down as increases from to . Thus, we select the moderate value of to reach a balance between attack utility and imperceptibility. For text perturbation, exhibits a more profound influence on the black-box attacks. In our experiments, we strictly set for invisibility. However, attackers can adjust the value of in accordance with their demands to trade off the attack efficacy and the perturbation stealthiness.
5 Conclusion
This paper delves into the challenging task of launching universal adversarial attacks against VLP models and proposes an effective solution that achieves superior performance using only one universal pair of image-text perturbations. We begin by revealing the unsatisfactory results of existing UAP methods and empirically explaining the underlying reasons. Based on our analysis, we propose to break the crucial cross-modal alignment in VLP models by designing a contrastive-learning generative UAP framework that leverages both unimodal and multimodal information to enhance the attacks. Extensive experiments validate the efficacy of the proposed algorithm on diverse VLP models and V+L tasks. We highlight that the proposed framework makes a significant step in exploring the classic universal adversarial attacks in VLP models and deepens our understanding of the mechanism of VLP models. We also hope that this paper can promote future research that explores more sophisticated defenses to strengthen the resilience of VLP models against adversarial attacks.
Appendix A Pseudocode of the Proposed Algorithm
We present the pseudocode of our proposed attack algorithm for image modality in Alg. 1. Note that the text attacks are completely symmetrical as illustrated in Sec. 3.3.
Appendix B Detailed Introductions to Datasets
Flickr30K (Plummer et al., 2015). Collected from the Flickr website, this dataset describes different items and activities, which becomes a standard benchmark for various V+L tasks. It contains 31,783 images, each of which has five associated captions. We use it for ITR tasks.
MSCOCO (Lin et al., 2014). The MSCOCO dataset is a rich and diverse dataset consisting of 123,287 images, each of which is annotated with approximately five sentences. We use this dataset to test the attack performance of ITR and IC tasks.
SNLI-VE (Xie et al., 2019). Originally proposed for natural language reasoning tasks, this dataset provides large-scale images and descriptions, where each image is annotated with several sentences and their logical relationship labels, including entailment, neutral, and contradiction. This dataset is used for VE tasks.
RefCOCO+ (Yu et al., 2016). An image dataset was selected from MSCOCO, which contains 19,992 images and 141,564 annotations. It is specially used for visual grounding (VG) tasks.
Appendix C More Training Details
We initialize the noise variable as a matrix. Meanwhile, the initial noise ’s dimensions in the text modality depend on the size of the hidden layer within the specific VLP model. Concretely, we set its dimension to for ALBEF, TCL, BLIP, and X-VLM, while for the CLIP model. When computing the multimodal contrastive loss , the temperature is set as 0.1. The generator is trained over 40 epochs with the Adam optimizer, utilizing a learning rate of . Following previous works Lu et al. (2023); Wang et al. (2024), we employ the attack success rate (ASR) as our quantitative measurement of our attack in ITR tasks by computing the extent the adversarial perturbations result in victim models’ performance deviations from the clean performance.
Appendix D Multimodal Alignment Destruction

### v3

@@ Appendix C More Experimental Results
| Flickr30K (R@5) | ALBEF | ETU | 68.54 | 77.68 | 14.41 | 14.43 | 4.1 | 6.41 | 6.11 | 16.97 | 11.57 | 23.11 | 9.36 | 13.3 |
| TCL | ETU | 14.43 | 16.23 | 84.88 | 80.43 | 2.4 | 5.8 | 5.6 | 17.37 | 13.27 | 24.27 | 7.24 | 10.98 | |
| X-VLM | ETU | 3.81 | 5.85 | 3.9 | 6.51 | 89.2 | 84.57 | 5.18 | 16.9 | 14.33 | 24.52 | 3.12 | 7.81 | |
| CLIP | ETU | 5.31 | 7.88 | 7.51 | 10.42 | 4.7 | 8.92 | 76.42 | 84.54 | 35.46 | 41.73 | 5.03 | 10.2 | |
| CLIP | ETU | 1.7 | 5.02 | 3.4 | 7.16 | 1.4 | 6.02 | 6.22 | 17.61 | 84.39 | 87.85 | 2.31 | 8.26 | |
| Ours | 4.41 | 19.8 | 6.41 | 25.19 | 4.8 | 23.49 | 18.76 | 43.82 | 90.34 | 88.12 | 8.95 | 26.89 | ||
| BLIP | ETU | 8.22 | 10.11 | 6.71 | 10.46 | 3.2 | 5.62 | 5.28 | 16.39 | 9.24 | 21.97 | 45.98 | 73.75 | |
| MSCOCO (R@5) | ALBEF | ETU | 81.73 | 88.76 | 13.45 | 11.51 | 9.01 | 9.32 | 16.85 | 20.74 | 22.6 | 28.86 | 10.96 | 12.02 |
| TCL | ETU | 44.94 | 36.45 | 90.21 | 84.54 | 13.46 | 13.1 | 19.52 | 23.03 | 30.81 | 36.37 | 25.63 | 25.11 | |
| X-VLM | ETU | 11.03 | 11.11 | 10.22 | 9.24 | 94.36 | 90.02 | 17.56 | 20.8 | 33.45 | 38.14 | 10.08 | 10.12 | |
| CLIP | ETU | 15.89 | 16.01 | 18.61 | 16.11 | 16.45 | 16.61 | 93.12 | 94.66 | 72.97 | 75.5 | 17.1 | 19.26 | |
| CLIP | ETU | 9.55 | 10.34 | 9.9 | 10.95 | 10.98 | 11.96 | 22.63 | 26.84 | 90.7 | 92.07 | 9.69 | 12.25 | |
| Ours | 16.83 | 31 | 19.86 | 34.54 | 18.84 | 34.02 | 50.21 | 59.2 | 90.94 | 90.43 | 25.5 | 44.89 | ||
| BLIP | ETU | 29.6 | 29.6 | 26.95 | 21.97 | 17.67 | 16.14 | 20.83 | 24.27 | 32.77 | 36.72 | 76.16 | 80.69 | |
| Flickr30K (R@10) | ALBEF | ETU | 65.8 | 74.89 | 10 | 9.36 | 2.8 | 3.79 | 2.43 | 11.04 | 6.95 | 15.79 | 6.62 | 8.57 |
| TCL | ETU | 11.4 | 11.35 | 82.2 | 77.33 | 1.6 | 3.38 | 3.04 | 10.63 | 7.06 | 17.14 | 4.91 | 7.34 | |
| X-VLM | ETU | 1.9 | 3.48 | 2.4 | 3.48 | 87.1 | 81.87 | 3.14 | 10.91 | 8.59 | 16.66 | 2.01 | 4.72 | |
| CLIP | ETU | 3.6 | 4.75 | 4.6 | 6.43 | 2.7 | 5.47 | 68.17 | 78.93 | 27.2 | 32.94 | 3.11 | 6.35 | |
| CLIP | ETU | 0.6 | 2.64 | 1.6 | 3.86 | 0.7 | 3.48 | 3.65 | 11.32 | 80.37 | 85.18 | 1.2 | 5.31 | |
| Ours | 2.4 | 14.59 | 3.5 | 18.36 | 2.3 | 17.95 | 11.75 | 34.23 | 86.4 | 83.83 | 5.52 | 20.86 | ||
| BLIP | ETU | 6.5 | 6.52 | 4.9 | 6.57 | 1.8 | 3.32 | 2.63 | 10.52 | 5.52 | 14.57 | 42.43 | 71.26 | |
| MSCOCO (R@10) | ALBEF | ETU | 81.14 | 87.58 | 9.08 | 7.92 | 5.44 | 6.33 | 12.62 | 16.15 | 17.71 | 23.05 | 7.63 | 9.15 |
| TCL | ETU | 38.6 | 30.39 | 88.56 | 82.6 | 8.92 | 9.3 | 14.61 | 18.43 | 25.24 | 30.06 | 20.37 | 21.41 | |
| X-VLM | ETU | 7.41 | 7.37 | 6.76 | 6.27 | 93.17 | 88.66 | 12.93 | 16.43 | 28.32 | 32.47 | 6.73 | 8.01 | |
| CLIP | ETU | 11.05 | 11.5 | 13.67 | 11.67 | 11.21 | 12.14 | 91.47 | 93.8 | 67.7 | 71.14 | 13.03 | 15.4 | |
| CLIP | ETU | 6.05 | 7.02 | 6.54 | 7.54 | 6.97 | 8.36 | 17.65 | 21.66 | 88.13 | 90.16 | 6.73 | 9.45 | |
| Ours | 10.77 | 24.11 | 13.34 | 27.53 | 12.31 | 28.03 | 41.33 | 52.02 | 88.28 | 87.14 | 20.26 | 39.43 | ||
| BLIP | ETU | 23.63 | 24.4 | 19.69 | 16.37 | 11.55 | 11.86 | 16.3 | 19.66 | 26.04 | 30.64 | 73.88 | 77.63 | |
@@
| 12.02 | 30.75 | 14.39 | 35.08 | 11.41 | 30.79 | 37.32 | 56.05 | 40.17 | 56.39 | 19.66 | 37.33 | |
| 57.55 | 67.4 | 37.06 | 49.45 | 10.7 | 28.48 | 37.49 | 58.3 | 40.87 | 58.39 | 23.33 | 39.44 | |
| 76.46 | 82.46 | 56.52 | 62.61 | 14.33 | 33.61 | 42.98 | 62.81 | 46.11 | 65.58 | 27.13 | 46.44 | |
| + | 81.09 | 83.71 | 48.76 | 56.54 | 17.58 | 35.72 | 41.5 | 64.72 | 47.41 | 70.34 | 35.96 | 51.76 |
| + | 65.20 | 72.71 | 36.13 | 50.06 | 18.63 | 36.74 | 42.23 | 65.17 | 50.91 | 69.78 | 36.91 | 50.69 |
| + | 90.13 | 88.82 | 62.11 | 64.48 | 20.53 | 39.38 | 43.1 | 65.93 | 54.4 | 72.51 | 44.79 | 56.36 |
Performance under Defenses. We next analyze several defense strategies to mitigate the potential harm from C-PGC. Concretely, we totally align with TMM [41] and consider several input preprocessing-based schemes, including image smoothing [10] (Gaussian, medium, average smoothing), JPEG compression [14], NRP [33], and the DiffPure [34], a powerful purification defense using diffusion models. For adversarial text correction, we choose the LanguageTool (LT) [41], which has been widely adopted in various scenarios due to its universality and effectiveness.
The attack results in Table 8 demonstrate that the proposed attack still attains great ASR against different powerful defenses. It also indicates that NRP+LT would be a decent choice to alleviate the threat brought by C-PGC. Another noteworthy finding is that, although DiffPure [34] exhibits remarkable performance in defending attacks in classification tasks, its ability is greatly reduced in V+L scenarios since the denoising process could also diminish some texture or semantic information that is critical for VLP models, thereby acquiring unsatisfactory defense effects.
Appendix D Rationality behind the Loss Design
It is widely acknowledged that contrastive learning serves as a powerful and foundational tool for modality alignment in VLP models, establishing a nearly point-to-point relationship between image and text features. Since contrastive learning can establish robust and precise alignment, leveraging the same technique to disrupt the established alignments is also promising to yield effective performance.
Taking image attack as an example, the underlying principle behind our contrastive learning-based attack can be understood from two perspectives:
Leverage the originally matched texts as negative samples to push aligned image-text pairs apart. This broadly corresponds to the objective of traditional untargeted adversarial attacks.
Additionally, the proposed paradigm introduces dissimilar texts as positive samples to further pull the adversarial image out of its original subspace and relocate it to an incorrect feature area.
By simultaneously harnessing the collaborative effects of push (negative samples) and pull (positive samples), the proposed contrastive framework achieves exceptional attack performance, which has been validated by comprehensive experimental results. Besides, we also explore several potential alternative loss functions that more directly align with the common goal of untargeted attack in Table 9, including maximizing the cosine distance or MSE distance between features of matched image-text pairs.
Recall that and denote our designed contrastive loss and unimodal loss terms respectively. As observed, the integration of consistently brings significant ASR improvements, verifying the rationality and superiority of the adopted contrastive loss.
| Method | ALBEF | TCL | CLIP | CLIP | ||||||||
| P | R | F1 | P | R | F1 | P | R | F1 | P | R | F1 | |
| Co-Attack [48] | 0.8328 | 0.8589 | 0.8455 | 0.8325 | 0.8588 | 0.8453 | 0.8269 | 0.8526 | 0.8394 | 0.8271 | 0.8530 | 0.8397 |
| SGA [29] | 0.8389 | 0.8654 | 0.8518 | 0.8376 | 0.8646 | 0.8509 | 0.8416 | 0.8697 | 0.8553 | 0.8378 | 0.8650 | 0.8511 |
| Ours | 0.8891 | 0.8613 | 0.8748 | 0.8924 | 0.8687 | 0.8802 | 0.8746 | 0.8684 | 0.8713 | 0.8948 | 0.8842 | 0.8893 |
Appendix E Semantic Similarity Analysis
The basic objective of untargeted adversarial attacks is to fool the victim model to output incorrect predictions [12], while the attacker is supposed to preserve semantic similarity between original and adversarial samples to ensure attack imperceptibility. In our implementation, we follow the rigorous setup in prior works [48, 29, 41] that modify only one single word to preserve the attack stealthiness. To quantitatively analyze the influence, we provide the BERT scores [50], which calculate the P (precision), R (recall), and F1 (F1 score), to measure the semantic distance between 5,000 clean and adversarial sentences in Table 10. Note that we provide existing well-acknowledged sample-specific algorithms Co-Attack [48] and SGA [29] as references.
As observed, C-PGC generally acquires higher similarity scores than existing sample-specific methods across various surrogate models, which validates that C-PGC achieves an eligible perturbation strategy in terms of text perturbation imperceptibility. Basically, the lower semantic similarity of sample-specific approaches stems from their word-selection mechanism, which maximizes the semantic distance tailored to every input sentence for attack enhancement. I.e., these methods select the adversarial word that maximizes the distance between the original and perturbed texts for every input sentence, which inherently leads to relatively larger semantic deviations. This highlights that our method achieves a better balance between efficacy and stealthiness.
Appendix F Multimodal Alignment Destruction